OpenStudy (anonymous):

Multiply and simplify: 10x(6x2 + 7x + 9)

OpenStudy (imstuck):

Are those "x"'s variables, or do you mean for them to be multiplication signs? Please clarify!

OpenStudy (solomonzelman):

\[10x(6x^2 + 7x + 9)~=~(10 \times 6x^2)+(10 \times7x)+(10 \times9)~=\]\[60x^2+70x+90~.\]

OpenStudy (montesanna99):

It should be 60x^3+70x^2+90x

OpenStudy (anonymous):

yeah, it should be ^3 and not ^2

OpenStudy (anonymous):

if you have the answer @Wolverine32 close the question